# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v2

@@ Abstract
Urban traffic regulation policies are increasingly used to address congestion, emissions, and accessibility in cities, yet their impacts are difficult to assess due to the socio-technical complexity of urban mobility systems. Recent advances in data availability and computational power enable new forms of model-driven, simulation-based decision support for transportation policy design. This paper proposes a novel simulation paradigm for the ex-ante evaluation of both direct impacts (e.g., traffic conditions, modal shift, emissions) and indirect impacts spanning transportation-related effects and economic accessibility. The approach integrates a multi-layer urban mobility model combining a physical layer of mobility flows and emissions with a social layer capturing behavioral responses and adaptation to policy changes. Real-world data are used to instantiate the current “as-is” scenario, while policy alternatives and behavioral assumptions are encoded as model parameters to generate multiple “what-if” scenarios. The framework supports systematic comparison across scenarios by analyzing variations in simulated outcomes induced by policy interventions. The proposed approach is illustrated through a case study that aims to assess the impacts of the introduction of broad urban traffic restriction schemes. Results demonstrate the framework’s ability to explore alternative regulatory designs and user responses, supporting informed and anticipatory evaluation of urban traffic policies.
@@ 1 Introduction
This paper contributes to this line of research by proposing a model-driven, simulation-based paradigm for decision support in urban mobility policy design. The objective is to enable the ex-ante evaluation of both direct impacts—such as traffic conditions, congestion, modal split, and emissions—and indirect impacts arising in related domains. These include transportation-related effects (e.g., parking displacement), social impacts (e.g., differential constraints on mobility for vulnerable groups), and economic consequences (e.g., accessibility of city centers, tourism, and local commerce). By supporting systematic exploration of policy alternatives before implementation, the proposed approach aims to reduce uncertainty and unintended consequences in urban traffic regulation.
@@
The approach is inherently simulation-based, allowing the representation of the current “as-is” state of the mobility system alongside multiple “what-if” scenarios corresponding to alternative traffic policies and structural changes in urban mobility. This enables controlled experimentation on policy options that would be impractical, costly, or ethically problematic to test in real-world settings.
@@
At its core, the framework relies on an urban mobility model that explicitly integrates a physical layer, including the mobility traffic flows and pollutant emissions, with a social layer capturing individual and collective behavior, adaptation mechanisms, and varying degrees of behavioral rigidity. This coupling is essential to represent the socio-technical nature of urban mobility systems and to account for behavioral responses to policy interventions.
The “as-is” scenario is instantiated using directly measured and estimated quantities derived from available data sources. These include mobility-related indicators such as vehicle volumes, origin–destination matrices, and emissions, represented at both macro and micro levels. Policy alternatives are then defined through model parameters that encode endogenous events (e.g., public authority regulations), exogenous events (e.g., infrastructure disruptions, changes in vehicle fleets), and assumptions on user responses and behavioral adaptation.
@@
Simulation outputs characterize the system behavior under each “what-if” scenario, with particular emphasis on variations in outcomes such as traffic patterns and emissions, induced by changes in model parameters. The framework also supports systematic comparison across scenarios and against the baseline “as-is” condition. We remark that the scope of this work is not causal inference or factor discovery: we assume that relevant factors and their relationships to the modeled phenomena are given and encoded in the model structure, and policy analysis is conducted by manipulating input parameters, generating alternative system states, and comparing their simulated outcomes.
@@
To illustrate the proposed approach, we apply it to a case study in the city of Bologna, Italy, a major urban center characterized by a large resident population and a complex, multimodal mobility system. The case study focuses on assessing the direct and indirect impacts of a hypothetical newly introduced traffic fee aimed at regulating vehicular circulation within a designated urban area. Building on the city’s existing limited traffic zone in the historical center, we simulate—for experimental purposes—an additional circulation restriction applied to a broader area. This setting enables us to demonstrate the ability of the proposed framework to construct, analyze, and compare multiple “what-if” scenarios, reflecting alternative regulatory designs and heterogeneous user responses, and to support informed evaluation of traffic regulation policies before their implementation.
@@ 2.1 Policy application and evaluation
In recent decades, transportation policies have assumed a central role in the international political agenda, driven by the need to balance the traffic externalities with economic growth, social equity, accessibility, and environmental objectives. The literature has identified various types of traffic policy interventions, reflecting different strategic approaches to address transport challenges. Operationally, these policies can be categorized according to whether they target the supply side or the demand side of the transport system.
Supply strategies are designed to increase the quality of the supply and, hence, its traffic capacity and efficiency. Examples of such traffic policies are the modifications to the transport infrastructure, both aimed at equipping the system with new facilities (e.g., the construction of a new highway or a new railway connection) and at enhancing the existing system through maintenance and expansion (e.g., the increase of the capacity of an existing road through the introduction of new lanes). Similarly, the technological enhancement of the supply, towards the constitution of Intelligent Transport Systems, is designed to monitor and manage traffic in real-time and provide information to travelers about traffic conditions, enabling better route choices and reducing congestion. Another key example of supply policy is the optimization of the public transport systems, where the scheduling and the routing are recalibrated to properly balance the mobility demand with the mobility offer. As demonstrated by Rhoads and Shogren (2006), supply-side policies tend to have a higher social discount rate and are often preferred to obtain immediate and tangible changes in the transport system. However, they may induce additional demand and generate the so-called induced traffic, sometimes offsetting long-term benefits (Goodwin, 1996; Downs, 2004; Hymel, 2019).
Alongside, demand-side traffic policies address the mitigation of congestion by altering the demand for transportation (Lindsey et al., 2023; Small et al., 2024). These demand-side approaches take various forms, including access regulations that limit entry and circulation in specific zones or roads, for all vehicles or specific categories. Taxation measures, such as fuel taxes and vehicle ownership taxes, also play a key role in discouraging the use and/or the possession of private vehicles, in favor of other transport modes. Similarly, adjustments to public transport fares can incentivize or disincentivize ridership, while carpooling initiatives may encourage shared vehicle use. Finally, pricing schemes introduce user charges for accessing certain roads or zones, often to limit vehicular traffic congestion.
@@
As widely described by de Palma and Lindsey (2011), the enactment of pricing policies serves to internalize the marginal social cost imposed by travelers, who contribute to collective traffic congestion and its negative externalities, and are therefore required to bear their share of this collective cost. Beyond the original theoretical framework proposed by Pigou (1920) for computing optimal traffic fees, practical implementations of pricing policies exhibit considerable variation. Common approaches include area-based charging (i.e., tolls for entering or passing through designated zones) and facility-based charging (i.e., tolls for using specific roads or lanes). These schemes also differ significantly in their temporal implementation, with some applying flat charges during the whole day and others during peak hours, while more advanced systems employ dynamic, real-time pricing strategies that adapt the pricing fee to current traffic conditions (Bonsall et al., 2007). Pricing schemes commonly differentiate by vehicle characteristics (e.g., fuel type, registration year, weight) and often include discounts or exemptions for specific vehicle categories and driver groups.
@@
Cost-Benefit Analysis (CBA) represents the standard tool for policy evaluation, being among the most widely discussed, tested, and used techniques (Koopmans and Mouter, 2020). In CBA, the societal value of a new policy is evaluated by comparing in monetary terms the total cost related to its implementation (e.g., planning and construction costs) and the benefits that it causes (e.g., savings in fuel consumption, savings in maintenance). While it is effective at assessing direct and tangible effects, CBA is often criticized for its difficulty in fully capturing externalities. Many of these are not easily linked to a clearly defined economic metric, which can result in distorted or biased estimates (Næss, 2020). Other comprehensive methodologies have emerged in response to the persistent gap between standard policy evaluation practices and the theoretically ideal comprehensive solutions (Oliveira and Pinho, 2010; Koopmans and Mouter, 2020; Næss, 2020). However, a major challenge remains: once the quantities of interest have been identified, one must propose a model that can capture the system’s response mechanisms to policy implementation and adequately adjust these quantities accordingly. Only by adopting such a modeling approach, outcome differences can be estimated rigorously and robustly, thereby allowing for their appropriate evaluation.
@@ 2.2 Socio-technical mobility systems: models and challenges
Since their initial development in the 1950s, the goal of mobility models has been to provide a systematic methodology for quantifying the effects of changes in the underlying transportation systems (van Nes and de Jong, 2020). For example, the four-step model has long been used to structure mobility analysis and to anticipate the economic impact of large-scale infrastructure projects (McNally, 2007). Similarly, System Dynamics offers an aggregate representation of system behaviour to explore how complex dynamics may unfold over time in strategic long-term policy scenarios (Shepherd, 2014). Recent methodological advances and new applications have further enabled researchers to address increasingly specific questions and to study diverse transportation contexts. Macroscopic models capture aggregate system behaviour (Wilson, 1970; McNally, 2007; Shepherd, 2014), while microscopic models represent individual agents and their interactions (Rilett, 2001; Horni et al., 2016; Lopez et al., 2018; PTV Group, 2025). Both are widely used to explore what-if scenarios, such as assessing how mobility patterns may change under alternative policy interventions, as illustrated in the works by Rieser et al. (2008); Brand et al. (2019); Melkonyan et al. (2020); Daniel et al. (2021); Asgarpour et al. (2023).
@@
While models differ in focus and detail, they address and integrate fundamental components to capture the essential elements that allow researchers to describe, understand, and analyze the complex phenomena under study. First, mobility systems are intrinsically socio-technical complex systems (Batty, 2008; McPhearson et al., 2016; Luca et al., 2024). Social components (e.g., people, organizations, cultures, processes) and techniques (e.g., technologies, infrastructures, tools) are deeply interconnected and influence each other. They integrate physical infrastructures, technologies, and digital systems with individual behaviors and social practices components, and the interactions of these closely connected components give rise to complex, emergent behaviors and trade-offs inherent to large-scale socio-technical systems (Sussman, 2000). Capturing this complexity is a key challenge and requires comprehensive, fair, and efficient models capable of supporting and improving informed decision-making processes on policy development and scenario analysis.
@@
Uncertainty is another inherent feature of mobility systems, closely related to their complexity, as it measures “any deviation from the unachievable ideal of completely deterministic knowledge of the relevant system” (Walker et al., 2003). Cities face a wide range of unpredictable events, from natural disasters and health emergencies to sudden economic shifts or policy changes. Even in stable conditions, human behavior adds another layer of unpredictability: people adjust their habits, react differently to incentives, and sometimes resist new regulations in ways that defy precise modeling (Avineri and Prashker, 2004). Uncertainty emerges across multiple dimensions. For instance, data limitations may introduce bias and incompleteness, while, at the modeling stage, simplifications and context-dependent assumptions can obscure complex causal relationships. Moreover, in forecasting, small inaccuracies tend to accumulate over time, potentially leading to large deviations between projected and observed outcomes (Mannucci et al., 2023).
@@ 3.1 Motivating case: traffic regulation pricing policy
The motivating case considers an urban area characterized by recurring congestion during peak hours and significant levels of atmospheric emissions. In such contexts, implementing traffic regulation policies is crucial, yet their evaluation proves challenging due to the complex interactions between user behavior, network characteristics, and regulatory parameters. The policy under study is a traffic pricing system, where tolls vary according to temporal and categorical criteria, with exemptions for specific user groups and a defined application schedule.
The complexity of the case becomes evident when observing how demand responds to the implemented measures. Users’ choices are heterogeneous and influenced by factors such as travel time, availability of alternative transport modes, and individual habits. The overall effect of the policy on traffic and emissions cannot be intuitively determined in a purely intuitive manner, as interactions between individual behaviors and regulatory measures generate nonlinear dynamic phenomena. In addition, uncertainty may be present in the unpredictable strategic response of users to the applied regulation. This clearly demonstrates the need for a systemic approach capable of simultaneously modeling the physical urban layer, user strategies, and policy parameters. Evaluating the effectiveness of interventions requires a framework that can represent both the policy and the behavioral diversity of users, while preserving the dynamic interdependencies within the system.
The main contribution of this framework is to generate comparable “as-is” and “what-if” scenarios, including their behavioural implications. This scenario generation constitutes the analytical starting point for policy appraisal: methods such as cost-benefit analysis depend precisely on such consistent “as-is” and counterfactual configurations, and the framework is designed to supply them in a transparent and reproducible manner. By examining these scenarios, the motivating case provides a concrete illustration of the challenges that the proposed framework aims to address. It clarifies why a systematic approach is necessary to link regulatory interventions, behavioral responses, and resulting system states in a coherent and reproducible manner. The case also lays the conceptual foundation for the framework’s components: scenarios representing the initial conditions of the system, policy parameters defining the intervention, and behavioral strategies representing user responses. Through this lens, the motivating case becomes a tangible demonstration of the complexity inherent in urban traffic systems and the rationale for adopting a structured modeling approach.
@@ 3.2 General framework architecture
Figure 1: Schematic representation of the architectural framework for “what-if” scenario analysis
As Figure 1 depicts, the model evaluates the current scenario and, together with a set of external control parameters, simulates the changes that lead to a modified scenario. This allows for the evaluation of the policy impacts by comparing the baseline and the modified configurations.
Within a spatio-temporal domain, a scenario is described by its core quantities of interest, which are the measurable variables that the model aims to consider, describe, and evaluate. They refer to the relevant aspects of reality, including physical and environmental features (such as the road network and the land use), social characteristics (such as the number of people with specific attributes), and the dynamics of the complex civic system. In the baseline scenario, the quantities of interest are inferred directly from the real-world data within the spatio-temporal domain of the analysis. Indeed, the baseline scenario represents the current conditions of the analysed framework and serves as a benchmark for measuring the impact of any intervention introduced. In contrast, the modified scenario represents the situation after the implementation of changes in the current system, and the values of the quantities of interest are estimated throughout the model.
The changes described above can result from either exogenous or endogenous events. Exogenous events originate from outside the system, such as changes in the environment, new regulations, or unexpected external shocks. Endogenous events, in contrast, emerge from within the system itself, arising from internal interactions or natural changes that develop over time within the system. In both cases, the goal is to model these events using a series of control parameters. Their definition is problem-dependent and has the objective of identifying the key characteristics of the event and schematizing them with a certain number of parameters.
@@
The development of the modified scenario begins from the original system conditions. These conditions are altered through the occurrence of an event, defined by control parameters, which directly modifies the system and indirectly the users’ habits in the system. The combined effects extend through the system, affecting the key variables under study. The impact evaluation is hence assessed by comparing the baseline initial measures with the new modified ones. The extent of the change is measured as the variation in the quantities of interest.
In this framework, variables can be defined either deterministically, with fixed values, or stochastically, as random variables following probability distributions. The probabilistic formulation allows us to explicitly represent input uncertainty and trace how it propagates through the model to the final estimates of the quantities of interest. By treating parameters probabilistically rather than as fixed values, the model can be evaluated repeatedly, with each simulation producing potentially different outputs that reflect the underlying uncertainty. An ensemble approach enables the quantification of output uncertainty, either through small variations in input values or, in the case of nondeterministic models, by using different initializations for the random processes.

### v3

@@ Abstract
Urban traffic regulation policies are increasingly used to address congestion, emissions, and accessibility in cities, yet their impacts are difficult to assess due to the socio-technical complexity of urban mobility systems. Recent advances in data availability and computational power enable new forms of model-driven, simulation-based decision support for transportation policy design. This paper proposes a novel simulation paradigm for the ex-ante evaluation of direct and indirect impacts, spanning traffic conditions, transportation-related effects and economic accessibility. The approach integrates a multi-layer urban mobility model combining a physical layer of mobility flows and emissions with a social layer capturing behavioral responses and adaptation to policy changes. Real-world data are used to instantiate the current as-is scenario, while policy alternatives and behavioral assumptions are encoded as model parameters to generate multiple what-if scenarios. The framework supports systematic comparison across scenarios by analyzing variations in simulated outcomes induced by policy interventions. The proposed approach is illustrated through a case study that aims to assess the impacts of the introduction of broad urban traffic restriction schemes. Results demonstrate the framework’s ability to explore alternative regulatory designs and user responses, supporting informed and anticipatory evaluation of urban traffic policies.
@@ 1 Introduction
This paper contributes to this line of research by proposing a model-driven, simulation-based paradigm for decision support in urban mobility policy design. The objective is to enable the ex-ante evaluation of direct impacts, such as traffic conditions, congestion, modal split, and emissions, and indirect impacts arising in connected domains. By supporting systematic exploration of policy alternatives before implementation, the proposed approach aims to reduce uncertainty and unintended consequences in urban traffic regulation.
@@
The approach is inherently simulation-based, allowing the representation of the current as-is state of the mobility system alongside multiple what-if scenarios corresponding to alternative traffic policies and structural changes in urban mobility. This enables controlled experimentation on policy options that would be impractical, costly, or ethically problematic to test in real-world settings.
@@
At its core, the framework relies on an urban mobility model that explicitly integrates a physical layer, including the mobility traffic flows and pollutant emissions, and a social layer capturing individual and collective behavior, adaptation mechanisms, and varying degrees of behavioral rigidity. This coupling is essential to represent the socio-technical nature of urban mobility systems and to account for behavioral responses to policy interventions.
@@
The as-is scenario is instantiated using directly measured and estimated quantities derived from available data sources. These include mobility-related indicators such as vehicle volumes, Origin–Destination matrices, and emissions, represented at both macro and micro levels. Policy alternatives are then defined through model parameters that encode endogenous events (e.g., public authority regulations), exogenous events (e.g., infrastructure disruptions, changes in vehicle fleets), and assumptions on user responses and behavioral adaptation. Simulation outputs characterize the system behavior under each of these what-if scenarios, with particular emphasis on variations in outcomes such as traffic patterns and emissions, induced by changes in model parameters. The framework also supports systematic comparison across scenarios and against the baseline as-is condition. We remark that the scope of this work is not causal inference or factor discovery: we assume that relevant factors and their relationships to the modeled phenomena are given and encoded in the model structure, and policy analysis is conducted by manipulating input parameters, generating alternative system states, and comparing their simulated outcomes.
@@
To illustrate the proposed approach, we apply it to a case study in the city of Bologna, Italy, a major urban center characterized by a large resident population and a complex, multimodal mobility system. The case study focuses on assessing the direct and indirect impacts of a hypothetical newly introduced traffic policy aimed at regulating vehicular circulation within a designated urban area. Building on the city’s existing limited traffic zone in the historical center, we simulate an additional circulation restriction applied to a broader area. This setting enables us to demonstrate the ability of the proposed framework to construct, analyze, and compare multiple what-if scenarios, reflecting alternative regulatory designs and heterogeneous user responses, and to support informed evaluation of traffic regulation policies before their implementation.
One key aspect of the proposed simulation approach is the attention given to uncertainty. This includes recognizing that variability in input data, modeling assumptions, and behavioral responses can influence both as-is and what-if scenario outcomes, and, where possible, quantifying its impact. While this aspect will not be the primary focus of the paper, the fact that the framework intrinsically supports it paves the way for future extensions toward fully uncertainty-aware evaluations.
@@ 2.1 Policy application and evaluation
In recent decades, transportation policies have assumed a central role in the international political agenda, driven by the need to balance the traffic externalities with economic growth, social equity, accessibility, and environmental objectives (Lindsey and Santos, 2020). The literature has identified various types of traffic policy interventions, reflecting different strategic approaches to address transport challenges (Santos et al., 2010a, b). Operationally, these policies can be categorized according to whether they target the supply side or the demand side of the transport system.
Supply strategies are designed to increase the quality of the supply and, hence, its traffic capacity and efficiency. Common examples of such strategies include upgrading infrastructure, implementing smart technology for real-time traffic updates, and refining public transport schedules to better match commuter needs. As demonstrated by Rhoads and Shogren (2006), supply-side policies tend to have a higher social discount rate and are often preferred to obtain immediate and tangible changes in the transport system. However, they may induce additional demand and generate the so-called induced traffic, sometimes offsetting long-term benefits (Goodwin, 1996; Downs, 2004; Hymel, 2019). Alongside, demand-side traffic policies address the mitigation of congestion by altering the demand for transportation (Lindsey et al., 2023; Small et al., 2024). These policies take many forms, for instance: restricting vehicle access in certain zones, using taxes to discourage car ownership, adjusting public transport fares, and implementing pricing schemes to reduce traffic congestion.
@@
As widely described by de Palma and Lindsey (2011), the enactment of pricing policies serves to internalize the marginal social cost imposed by travelers: who contributes to collective traffic congestion and its negative externalities is required to bear their share of the collective cost. Beyond the original theoretical framework for computing optimal traffic fees (Pigou, 1920), practical implementations of pricing policies exhibit considerable variation. Common approaches include area-based charging (i.e., tolls for entering or passing through designated zones) and facility-based charging (i.e., tolls for using specific roads or lanes), with fees that vary from flat daily rates to advanced real-time pricing (Bonsall et al., 2007). Pricing schemes commonly differ by vehicle characteristics (e.g., fuel type, registration year, weight) and may include discounts or exemptions for specific vehicle categories and driver groups.
@@
Cost-Benefit Analysis (CBA) represents the standard tool for ex-ante policy evaluation, being among the most widely discussed, tested, and used techniques (Koopmans and Mouter, 2020). In CBA, the societal value of a new policy is evaluated by comparing in monetary terms the total cost related to its implementation (e.g., planning and construction costs) and the benefits that it causes (e.g., savings in fuel consumption, savings in maintenance). While it is effective at assessing direct and tangible effects, CBA is often criticized for its difficulty in fully capturing externalities, which are not easily linked to a clearly defined economic metric and can result in distorted or biased estimates (Næss, 2020). Other comprehensive methodologies have emerged in response to the persistent gap between standard policy evaluation practices and the theoretically solutions (Oliveira and Pinho, 2010; Koopmans and Mouter, 2020; Næss, 2020). However, a major challenge remains: once the quantities of interest have been identified, one must propose models that can capture the system’s response mechanisms to policy implementation and adequately adjust these quantities accordingly. The adoption of such modeling approaches is indeeed necessary to estimate outcome differences rigorously and robustly, thereby allowing for their appropriate evaluation.
@@ 2.2 Socio-technical mobility systems: models and challenges
Since their initial development in the 1950s, the goal of mobility models has been to provide a systematic methodology for quantifying the effects of changes in the underlying transportation systems (van Nes and de Jong, 2020). For instance, the four-step model is frequently used to forecast the impact of major infrastructure projects (McNally, 2007), while System Dynamics serves as a sample approach for exploring how complex long-term policies might evolve (Shepherd, 2014). Recent advances and new applications now enable researchers to address more specific questions and study diverse transportation contexts. Macroscopic models capture aggregate system behavior (Wilson, 1970; McNally, 2007; Shepherd, 2014), while microscopic models represent individual agents and their interactions (Rilett, 2001; Horni et al., 2016; Lopez et al., 2018; PTV Group, 2025), and both are widely used to explore what-if scenarios, such as assessing how mobility patterns may change under alternative policy interventions (Rieser et al., 2008; Brand et al., 2019; Melkonyan et al., 2020; Daniel et al., 2021; Asgarpour et al., 2023).
@@
While models differ in focus and detail, they address and integrate fundamental components to capture the essential elements for describing, understanding, and analyzing the complex phenomena under study. First, mobility systems are intrinsically socio-technical complex systems (Batty, 2008; McPhearson et al., 2016; Luca et al., 2024). Social components (e.g., people, organizations, cultures, processes) and techniques (e.g., technologies, infrastructures, tools) are deeply interconnected and influence each other. They integrate physical infrastructures, technologies, and digital systems with individual behaviors and social practices components, and the interactions of these closely connected components give rise to complex, emergent behaviors and trade-offs inherent to large-scale socio-technical systems (Sussman, 2000). Capturing this complexity is a key challenge and requires comprehensive, fair, and efficient models capable of supporting and improving informed decision-making processes on policy development and scenario analysis.
@@
Uncertainty is another inherent feature of mobility systems, closely related to their complexity, as it measures “any deviation from the unachievable ideal of completely deterministic knowledge of the relevant system” (Walker et al., 2003). Uncertainty emerges across multiple dimensions: data limitations may introduce bias and incompleteness; simplifications and context-dependent assumptions at the modeling stage can obscure complex causal relationships; small inaccuracies in forecasting accumulate over time, potentially leading to large deviations between projected and observed outcomes (Mannucci et al., 2023). Cities face a wide range of unpredictable events, from natural disasters and health emergencies to sudden economic shifts or policy changes. Even in stable conditions, human behavior adds another layer of unpredictability: people adjust their habits, react differently to incentives, and sometimes resist new regulations in ways that defy precise modeling (Avineri and Prashker, 2004).
@@ 3.1 Motivating case: traffic regulation pricing policy
The illustrative case driving this study considers an urban area characterized by recurring congestion during peak hours and significant levels of atmospheric emissions. In such contexts, implementing traffic regulation policies is crucial, yet their evaluation proves challenging due to the complex interactions between user behavior, network characteristics, and regulatory parameters.
In the scope of this paper, the policy under study focuses on a specific regulatory action, namely, the introduction of a traffic pricing system where tolls vary according to temporal and categorical criteria, with exemptions for specific user groups and a defined application schedule. Even with this limited focus, the complexity of the case becomes evident when observing how demand responds to the implemented measures. Users’ choices are heterogeneous and influenced by factors such as travel time, availability of alternative transport modes, and individual habits. The overall effect of the policy on traffic and emissions cannot be determined in a purely intuitive manner, as interactions between individual behaviors and regulatory measures generate nonlinear dynamic phenomena. In addition, uncertainty may be present in the unpredictable strategic response of users to the applied regulation. This clearly demonstrates the need for a systemic approach capable of simultaneously modeling physical urban layers, user strategies, and policy parameters. Evaluating the effectiveness of interventions requires a framework that can represent both the policy and the behavioral diversity of users, while preserving the dynamic interdependencies within the system.
The main contribution of this framework is to generate comparable as-is and what-if scenarios, and to provide the analytical baseline for policy appraisal. By producing consistent as-is and what-if scenarios, it ensures that ex-post evaluations are grounded with explicit assumptions and modeling choices, and reproducible methods and procedures. The general framework is introduced in Section 3.2, while Section 3.3 introduces a motivating case: it provides a concrete illustration of the challenges which the proposed framework aims to address, and clarifies why a systematic approach is necessary to link regulatory interventions, behavioral responses, and resulting system states in a coherent and reproducible manner.
@@ 3.2 General framework architecture
As-is Scenario Model Control and Behavioral ParametersWhat-if Scenario Impact Evaluation
Figure 1: Schematic representation of the architectural framework for what-if scenario analysis.
Within a spatio-temporal domain, a scenario is described by its core quantities of interest, which are the measurable variables that the model aims to consider, describe, and evaluate. They refer to the relevant aspects of reality, including physical and environmental features (such as the road network and the land use), social characteristics (such as the number of people with specific attributes), and the dynamics of the complex civic system. In the as-is scenario, the quantities of interest are inferred directly from the real-world data within the spatio-temporal domain of the analysis. Indeed, the as-is scenario represents the current conditions of the analysed framework and serves as a benchmark for measuring the impact of any intervention introduced. In contrast, the what-if scenario represents the situation after the implementation of changes in the current system, and the values of the quantities of interest are estimated throughout the model.
The changes described above can result from either endogenous or exogenous events. Endogenous events emerge from within the system itself, arising from internal interactions or natural changes that develop over time within the system. Exogenous events originate from outside the system, either being unforeseen (e.g., unexpected changes in the environment or external shocks) or being imposed by external agents (e.g., new regulations). In both cases, the goal is to model these events using a series of control parameters. Their definition is problem-dependent and has the objective of identifying the key characteristics of the event and schematizing them with a certain number of parameters.
@@
The development of the what-if scenario begins from the original system conditions, as illustrated in Figure 1. These conditions are altered through the occurrence of an event, defined by control parameters, which directly modifies the system and indirectly the users’ habits in the system. The combined effects extend through the system, affecting the key variables under study. The impact evaluation is hence assessed by comparing the as-is initial measures with the new what-if ones. The extent of the change is measured as the variation in the quantities of interest.
In this framework, variables can be defined either deterministically, with fixed values, or stochastically, as random variables following probability distributions. The probabilistic formulation allows us to explicitly represent input uncertainty and trace how it propagates through the model to the final estimates of the quantities of interest. By treating parameters probabilistically rather than as fixed values, the model can be evaluated repeatedly, with each simulation producing potentially different outputs that reflect the underlying uncertainty. An ensemble approach enables the quantification of output variability, transforming single-point forecasts into a range of probable outcomes. This allows for the derivation of confidence intervals that provide a more statistically-rigorous foundation for comparing the as-is and what-if scenarios.